OpenStudy (anonymous):

Combine as indicated by the signs. 4+a/a-3 + 4/5

OpenStudy (anonymous):

2.8 or 14/5

OpenStudy (anonymous):

20 + 5a + 4a -12 ---------------- 5(a - 3) 9a +8 -------- 5(a - 3)

OpenStudy (anonymous):

lcd is 5(a-3) = 20 + 5a + 4a - 12 ------------------- 5(a-3) = 9a + 8 -------- 5(a-3) and i'll have to learn to type faster - lol!

OpenStudy (anonymous):

lol your fine! :)

OpenStudy (anonymous):

ty both